in a large restaurant an average of 3 out of 5 customers will ask for water with their waiter. a random sample of 50 customers is randomly selected. what is the probability that less than 34 ask for water with their meals?
Answer #1

P(Ask for water) = 3/5 = 0.60

Mean = np = 50*0.60 = 30

Standard deviation = = 3.464

Hence by using normal approximation to binomial:

P(Less than 34 asks for water)

= P(X < 33.5) [Continuity Correction]

= P(z < (33.5 - 30)/3.464)

= P(z < 1.01)

= 0.8438